2x/3= -9 Im not sure what to do here?
Answer by checkley79(3341)

2x/3=-9 cross multiply.
2x=-9*3
2x=-27
x=-27/2
x=-13/5 ans.
Proof:
2*-13.5/3=-9
-27/3=-9
-9=-9